# Protap: A Benchmark for Protein Modeling on Realistic Downstream Applications

### v2

@@ 3.3 Applications and Domain Models
In this subsection, we introduce the applications adopted in Protap, categorized into specialized and general. Specialized applications focus on specific proteins or biological processes, while general applications broadly apply across diverse proteins. We briefly describe each application along with representative domain models. A complete list of implemented domain models is provided in Tab. 2, and detailed experimental setups including datasets and metrics are available in Appendix C.
Specialized Application 1: Enzyme-Catalyzed Protein Cleavage Site Prediction (PCS). As illustrated in Fig. 1(b), proteolytic enzymes will first recognize specific amino acid sequences or structural motifs within substrate proteins with the enzyme’s active sites. Then, the enzymes catalyze the cleavage of peptide bonds at the cleavage site. This fundamental biological process regulates protein activity, turnover, and signaling across diverse biological contexts. Predicting the protein cleavage sites under the catalysis of enzymes has many crucial applications, such as enzyme engineering and therapeutic target identification. For example, in the design of enzyme inhibitors or prodrugs, identifying key cleavage peptides under the catalysis of the HIV enzyme could enhance drug specificity [46]. The enzyme-catalyzed protein cleavage site prediction can be formulated as a residue-level binary classification task. Specifically, let $\mathcal{P}^{s}$ and $\mathcal{P}^{e}$ denote the protein substrate and enzyme. The cleavage site predictor aims to learn the following function:
$f:(\mathcal{P}^{s},\mathcal{P}^{e})\rightarrow\{0,1\}^{|\mathcal{P}^{s}|}.$ (1)
The majority of existing methods, such as Procleave [47], construct enzyme-specific features to identify the cleavage sites under the enzyme of interest. Recently, UniZyme [22] has leveraged both enzyme and substrate encoders to build a unified cleavage site predictor that generalizes across various enzymes.
Specialized Application 2: Targeted Protein Degradation by Proteolysis-Targeting Chimeras (PROTACs). A Proteolysis-targeting chimeras (PROTAC) is a heterobifunctional molecule consisting of three components: a ligand for the targeted protein (commonly referred to as the warhead), a chemical linker, and a ligand that recruits an E3 ubiquitin ligase. PROTACs have emerged as powerful tools for selectively degrading disease-associated proteins via the ubiquitin-proteasome system [48]. As illustrated in Fig. 1(b), the PROTAC-mediated degradation process begins with simultaneous binding of the PROTAC to the target protein and an E3 ligase, resulting in a ternary complex. This complex subsequently facilitates the transfer of ubiquitin from the E2 enzyme to the target protein, ultimately leading to its proteasomal degradation. PROTACs offer several advantages over conventional treatments. First, due to their catalytic mode of action, PROTACs remain effective at lower doses, potentially reducing side effects [49, 50]. Second, unlike traditional drugs that rely on accessible binding pockets, PROTACs can target proteins previously considered undruggable and overcome resistance caused by mutations near active sites [51, 52]. However, the flexible ternary structure of PROTACs also introduces a more complex structure-activity relationship, making modeling this process more challenging. Formally, the prediction of PROTAC-mediated degradation of a target protein $\mathcal{P}^{t}$ by an E3 ligase can be defined as:
$f:(\underbrace{\text{Warhead},~{}\text{Linker},~{}\text{E3 ligand}}_{\text{PROTAC}},~{}\text{E3 ligase},~{}\mathcal{P}^{t})\rightarrow\{0,1\}$ (2)
To achieve the above task, initial efforts have been conducted to model the ternary complex formation induced by PROTAC molecules [23, 24]: DeepPROTACs employs GCNs to encode the ternary complex. ET-PROTAC considers the cross-talk between the PROTAC, target protein, and E3 ligase, which enables the modeling of the complex process in the targeted protein degradation.
General Application 3: Protein–Ligand Interactions (PLI). As illustrated in Fig. 1(b), protein–ligand binding refers to the highly specific and affinity-driven interaction between a protein and a small molecule ligand, resulting in a stable complex that can block the binding of other molecules [53]. The binding affinity can be quantified by the Gibbs free energy change $\Delta G$ of the protein-ligand complex, where a more negative $\Delta G$ indicates stronger binding affinity. Accurate prediction of protein–ligand binding affinity enables efficient identification of promising drug candidates and accelerates the optimization of therapeutic molecules. For example, kinases play critical roles in regulating cell growth and survival pathways implicated in cancer. Thus, predicting kinase–ligand binding affinity facilitates the effective screening of kinase inhibitors, which can selectively block these pathways. The screened kinase inhibitors are promising candidates for targeted cancer therapy. Formally, predicting protein–ligand binding affinity can be formulated as a regression task: $f:(\mathcal{P},\mathcal{M})\rightarrow\mathbb{R}$, where $\mathcal{P}$ denotes the protein and $\mathcal{M}$ denotes the molecule ligand. Some domain-specific methods, such as MONN [44], do not rely on structural information. To mitigate the false positive issue, KDBNet [43] applied an uncertainty recalibration technique to refine the uncertainty estimates.
General Application 4: Protein Function Annotation Prediction (PFA). To systematically represent protein functions, the Gene Ontology (GO) annotation framework was developed [54], providing structured annotations across Molecular Function (MF), Cellular Component (CC), and Biological Process (BP). For instance, in the Molecular Function (MF) aspect, a protein may be annotated with specific biochemical activities, such as immune Response or signal transduction. Among the known protein sequences, fewer than 1% have experimentally validated functional annotations. Predicting the GO terms of proteins helps researchers better understand protein functions and potentially guide the discovery and design of proteins with desired functional properties. Protein function prediction is inherently a multi-label classification task. Early methods relied on sequence alignment or direct modeling of the sequence. More recent approaches have progressively integrated additional modalities. For example, DeepFRI [45] combines protein structure and pre-trained sequence embeddings using a GCN, while DPFunc [21] further incorporates domain-level information.
General Application 5: Mutation Effect Prediction for Protein Optimization (MTP). Protein mutations refer to the substitution, insertion, or deletion of amino acids within protein sequences. As illustrated in Fig. 1(b), mutations are frequently associated with changes in functional properties such as stability, binding affinity, and pathogenicity. Computational models can explore the protein fitness landscape to map sequences or structures to functional properties, which can enable protein optimization with mutations. For example, mutations in antibodies or protein complexes can be systematically explored to discover variants with enhanced binding affinity [55]. Mutation effect prediction can be formulated as estimating changes in target properties (e.g., stability changes measured by $\Delta\Delta G$) resulting from protein mutations. Due to limited experimental annotations, mutation effect prediction typically relies on a zero-shot learning setting. Recently, protein language models (PLMs) have shown promise in mutation effect prediction due to their ability to capture patterns learned from evolutionary data. PLMs tend to assign higher probabilities to mutations that are consistent with evolutionary patterns, and these mutations are more likely to yield beneficial functional effects. Details of the zero-shot mutation effective prediction with PLMs can be found in Appendix C.5.
4 Experiments
In this section, we conduct empirical studies to compare domain models and various pretraining frameworks and strategies on the downstream applications described in Sec. 3.3.
4.1 Experimental Setup
Pre-training Dataset. For pre-training, we collect 542,378 protein structures from AlphaFold Protein Structure Database (AFDB) and obtain the corresponding amino acid sequences and protein family labels from UniProt [56, 42]. This curated dataset spans diverse and non-redundant protein structures, enabling a comprehensive evaluation of different training strategies. Additional dataset details are provided in Appendix E.2.
Table 3: Summary of datasets and metrics for various prediction tasks.
Application Category Data Source #Train #Test Metric
Pretraining Swiss-Prot [42] 542,378
@@ 4.1 Experimental Setup
Protein Cleavage Site Prediction Specialized MEROPS [57] 375 92 AUC, AUPR
@@
Targeted Protein Degradation Specialized PROTAC-DB 3.0 [58] 843 209 Acc, AUC
Protein–Ligand Interactions General DAVIS-KDBNet [43] 11,520 2,880 MSE, Pearson
Protein Function Annotation Prediction General DeepFRI [45], ProteinWorkshop [27] 23,760 2,023 Fmax, AUPR
Mutation Effect Prediction General ProteinGym [28] 2.4M AUC, Pearson
Downstream Application Dataset. Each downstream task uses the same dataset for both training-from-scratch and fine-tuning to ensure fair comparison. All datasets are collected from peer-reviewed sources and standardized through quality control and formatting, as detailed in Appendix C. Specifically, PROTACs data are from PROTAC-DB [58], cleavage data from MEROPS [57], protein–ligand interactions from KDBNet, function prediction from DeepFRI [45] (with structural data from ProteinWorkshop [27]), and mutation prediction from ProteinGym [28].
Downstream Task Training. Pretrained models serve as protein encoders in two settings: (i) Training from scratch, where all parameters are randomly initialized and trained end-to-end; (ii) Fine-tuning, where encoders are initialized with pretrained weights and frozen, while only task-specific heads are updated. All domain-specific models are trained from scratch.
Molecular Encoders in Downstream Applications. The applications of PLI and PROTACs involve various chemical molecules in the interactions with proteins, requiring a molecular encoder. To maintain consistency across interaction tasks, molecular components, drugs in PLI, warheads, linkers, and E3-ligands in PROTACs are encoded using randomly initialized GVP encoders. For domain-specific models, molecular encoders follow their original implementations.
Model Setups. We use fixed random seeds ${42,128,256,512,1024}$ for reproducibility, and report mean $\pm$ standard deviation over five runs. All models are trained with the Adam optimizer. The learning rate follows a linear warm-up over 10% of training steps, then cosine annealing to 1e-6, with a peak of 1e-4. More training details are provided in Appendix C.
@@
Table 4: Performance comparison across model architectures under different training strategies. The first line for each model (e.g., EGNN) denotes a randomly initialized protein encoder trained purely with downstream task supervision. The subsequent lines (e.g., w/MLM) represent pretrained encoders with frozen weights, where only the task-specific head is fine-tuned. Mutation effection prediction (MTP) is only applicable to the masked language modeling.
Model PCS PROTACs PLI PFA MTP
@@
D-Trans 97.60 $\pm$ 0.08 61.42 $\pm$ 1.30 88.28 $\pm$ 0.79 22.82 $\pm$ 2.95 80.08 $\pm$ 1.08 86.58 $\pm$ 0.28 0.416 $\pm$ 1.02 60.92 $\pm$ 0.22 19.57 $\pm$ 0.80 11.16 $\pm$ 0.67
w/MLM 96.53 $\pm$ 0.52 20.12 $\pm$ 2.86 84.10 $\pm$ 0.66 11.73 $\pm$ 1.41 77.78 $\pm$ 1.93 85.16 $\pm$ 0.89 0.494 $\pm$ 1.11 55.79 $\pm$ 0.50 17.92 $\pm$ 0.14 9.85 $\pm$ 0.09 -0.09 56.84
w/MVCL 95.23 $\pm$ 0.46 41.26 $\pm$ 10.3 83.50 $\pm$ 0.79 15.85 $\pm$ 1.30 74.16 $\pm$ 1.44 82.32 $\pm$ 0.81 0.550 $\pm$ 2.62 54.21 $\pm$ 0.08 15.87 $\pm$ 0.26 8.00 $\pm$ 0.20
w/PFP 96.40 $\pm$ 0.59 39.56 $\pm$ 5.55 85.51 $\pm$ 0.42 12.29 $\pm$ 1.26 74.76 $\pm$ 1.08 82.34 $\pm$ 0.09 0.445 $\pm$ 1.06 56.46 $\pm$ 0.01 17.46 $\pm$ 0.18 9.46 $\pm$ 0.12
ESM 97.23 $\pm$ 0.06 41.22 $\pm$ 0.25 86.34 $\pm$ 0.12 6.16 $\pm$ 0.07 78.46 $\pm$ 3.03 85.74 $\pm$ 3.47 0.491 $\pm$ 4.82 53.25 $\pm$ 1.26 49.79 $\pm$ 0.03 43.44 $\pm$ 0.17 43.05 73.48
4.2 Impacts of Pretraining Strategies and Architectures on Downstream Performance
We examine three pretraining tasks employed in Protap as described in Sec. 3.2, which are Masked Language Modeling (MLM), Multi-View Contrastive Learning (MVCL), and Protein Family Prediction (PFP). We pre-train EGNN, SE(3) Transformer (SE(3) Trans), GVP, ProteiBERT (ProtBERT), and D-Transformer (D-Trans) with all three tasks with 542k proteins to obtain pretraining encoders, which are listed in Tab. 2 and detailed introduced in Appendix E. For ESM, we directly utilize the publicly available pretrained weights, which were obtained via MLM training on 70M proteins. During the finetuning in the downstream task, their pretrained encoders are frozen for comparisons among strategies. For reference, we also train encoders with the downstream training set from scratch as the supervised encoder. The results are in Tab. 4, and we answer the following research questions.
RQ1: Can frozen pretraining encoders outperform a supervised encoder that trained on the downstream training set from scratch?
Overall, supervised encoders trained from scratch tend to outperform the pretraining encoders even when they are trained with large-scale datasets. For example, though ESM (650M parameters) already achieves competitive performance, ESM is often worse than the EGNN (10M parameters) encoder trained for the downstream tasks from the scratch. This suggests a degree of mismatch between the pretraining objectives and the downstream tasks, indicating that training from scratch enables the model to learn task-specific representations that are more aligned with the requirements of the downstream task.
RQ2: In what ways do variations in pretraining objectives impact the effectiveness of models on downstream tasks?
No significant patterns were observed regarding the impact of different pretraining strategies on downstream tasks. Various pretraining approaches do not exhibit a clear preference for specific downstream tasks. However, excluding models trained from scratch, those pretrained with PFP consistently achieve the best results on the PFA task compared to other pretraining methods, supporting the claim that proteins within the same family often share similar biological functions, and incorporating protein family information during pretraining aids functional prediction.
RQ3: Are models incorporating structures better than those purely using protein sequences?
Models that incorporate structural information generally achieve superior performance. Except for the PFA task, models such as EGNN, SE(3)-Trans, and D-Trans, which integrate spatial features, consistently outperform larger pretrained language models such as ESM and ProtBERT. This observation highlights two key insights. (i) The inclusion of three-dimensional structural information provides essential inductive biases that are absent in sequence-only architectures. (ii) Architectural alignment with biochemical properties can be more important than the sheer size of pretrained models in achieving strong performance on downstream protein-related tasks.
Table 5: Comparisons between general and domain models on protein cleavage site prediction.
General Architecture Domain Model
Metrics EGNN SE3 GVP ProtBERT D-Transformer Clipzyme Unizyme
C14.005 AUC(%$\uparrow$ 95.51 $\pm$ 0.03 81.76 $\pm$ 3.11 95.55 $\pm$ 0.46 95.49 $\pm$ 0.09 97.60 $\pm$ 0.08 92.27 $\pm$ 0.40 96.23 $\pm$ 0.10
AUPR(%$\uparrow$ 35.44 $\pm$ 0.06 02.07 $\pm$ 0.81 17.99 $\pm$ 5.02 40.49 $\pm$ 0.25 61.42 $\pm$ 1.30 43.23 $\pm$ 1.00 52.20 $\pm$ 0.90
M10.003 AUC(%$\uparrow$ 90.24 $\pm$ 0.32 67.12 $\pm$ 0.05 86.05 $\pm$ 1.68 89.83 $\pm$ 0.05 88.28 $\pm$ 0.79 82.50 $\pm$ 0.30 87.04 $\pm$ 0.20
AUPR(%$\uparrow$ 15.04 $\pm$ 1.18 01.01 $\pm$ 0.01 05.40 $\pm$ 0.52 10.74 $\pm$ 0.12 22.82 $\pm$ 2.95 05.81 $\pm$ 0.20 07.28 $\pm$ 0.40
Table 6: Comparisons between general and domain models on PROTACs.
General Architecture Domain Model
Metrics EGNN SE3 w/MLM GVP w/MVCL ProtBERT w/PFP D-Transformer w/PFP DeepPROTACs ET-PROTACs
Acc(%$\uparrow$ 80.19 $\pm$ 1.23 80.56 $\pm$ 2.08 70.75 $\pm$ 2.15 80.51 $\pm$ 1.03 80.08 $\pm$ 1.08 70.12 $\pm$ 0.83 78.87 $\pm$ 0.88
AUC(%$\uparrow$ 88.55 $\pm$ 1.78 87.64 $\pm$ 1.45 78.45 $\pm$ 1.30 87.46 $\pm$ 1.71 86.58 $\pm$ 0.28 77.59 $\pm$ 1.22 78.64 $\pm$ 1.21
Table 7: Comparisons between general and domain models on Protein-Ligand Interaction.
General Architecture Domain Model
Metrics EGNN SE3 GVP w/MVCL ProtBERT w/MVCL D-Transformer KDBNet MONN

### v3

@@ 4. Applications and Domain Models
In this subsection, we introduce the applications adopted in Protap, categorized into specialized and general. Specialized applications focus on interaction or biological processes, which include Enzyme-Catalyzed Protein Cleavage Site Prediction, Targeted Protein Degradation by Proteolysis-Targeting Chimeras and Protein–Ligand Interactions. General applications broadly apply across diverse proteins, which cover Protein Function Annotation Prediction and Mutation Effect Prediction for Protein Optimization. A complete list of implemented domain models is provided in Tab. 1, and detailed experimental setups including datasets and metrics are available in Appendix F.
@@
Specialized Application 1: Enzyme-Catalyzed Protein Cleavage Site Prediction (PCS).
Application Description of PCS. As illustrated in Fig. 3, proteolytic enzymes will first recognize specific amino acid sequences or structural motifs within substrate proteins with the enzyme’s active sites. Then, the enzymes catalyze the cleavage of peptide bonds at the cleavage site. This fundamental biological process regulates protein activity, turnover, and signaling across diverse biological contexts. Predicting the protein cleavage sites under the catalysis of enzymes has many crucial applications, such as enzyme engineering and therapeutic target identification. For example, in the design of enzyme inhibitors or prodrugs, identifying key cleavage peptides under the catalysis of the HIV enzyme could enhance drug specificity (Lv et al., 2015). The enzyme-catalyzed protein cleavage site prediction can be formulated as a residue-level binary classification task. Specifically, let $\mathcal{P}^{s}$ and $\mathcal{P}^{e}$ denote the protein substrate and enzyme. The protein cleavage site predictor aims to learn the following function:
(1) $f:(\mathcal{P}^{s},\mathcal{P}^{e})\rightarrow\{0,1\}^{|\mathcal{P}^{s}|}.$
PCS Domain Models. The majority of existing methods, such as Procleave (Li et al., 2020a), construct enzyme-specific features to identify the cleavage sites under the enzyme of interest. Recently, UniZyme (Li et al., 2025) has leveraged both enzyme and substrate encoders to build a unified cleavage site predictor that generalizes across various enzymes.
@@
Specialized Application 2: Targeted Protein Degradation by Proteolysis-Targeting Chimeras (PROTACs).
Application Description of PROTAC. A Proteolysis-targeting chimeras (PROTAC) is a heterobifunctional molecule consisting of three components: a ligand for the targeted protein (commonly referred to as the warhead), a chemical linker, and a ligand that recruits an E3 ubiquitin ligase. PROTACs have emerged as powerful tools for selectively degrading disease-associated proteins via the ubiquitin-proteasome system (Marei et al., 2022). As shown in Fig. 3, the PROTAC-mediated degradation process begins with simultaneous binding of the PROTAC to the target protein and an E3 ligase, resulting in a ternary complex. This complex subsequently facilitates the transfer of ubiquitin from the E2 enzyme to the target protein, ultimately leading to its proteasomal degradation. PROTACs offer several advantages over conventional treatments. First, due to their catalytic mode of action, PROTACs remain effective at lower doses, potentially reducing side effects (Neklesa et al., 2017; Toure and Crews, 2016). Second, unlike traditional drugs that rely on accessible binding pockets, PROTACs can target proteins previously considered undruggable and overcome resistance caused by mutations near active sites (Garber, 2022; Edmondson et al., 2019). However, the flexible ternary structure of PROTACs introduces a more complex structure-activity relationship, making modeling this process more challenging. The prediction of PROTAC-mediated degradation of a target protein $\mathcal{P}^{t}$ by an E3 ligase can be formulated as:
(2) $f:(\underbrace{\text{Warhead},~\text{Linker},~\text{E3 ligand}}_{\text{PROTAC}},~\text{E3 ligase},~\mathcal{P}^{t})\rightarrow\{0,1\}$
PROTACs Domain Models. Some initial efforts have been conducted to model the ternary complex formation induced by PROTAC molecules (Li et al., 2022; Cai et al., 2025): DeepPROTACs employs GCNs to encode the ternary complex. ET-PROTAC considers the cross-talk between the PROTAC, target protein, and E3 ligase, which enables the modeling of the complex process in the targeted protein degradation.
Specialized Application 3: Protein–Ligand Interactions (PLI).
Application Description of PLI. As illustrated in Fig. 3, protein–ligand binding refers to the highly specific and affinity-driven interaction between a protein and a small molecule ligand, resulting in a stable complex that can block the binding of other molecules (Miller and Dill, 1997). The binding affinity can be quantified by the Gibbs free energy change $\Delta G$ of the protein-ligand complex, where a more negative $\Delta G$ indicates stronger binding affinity. Accurate prediction of protein–ligand binding affinity enables efficient identification of promising drug candidates and accelerates the optimization of therapeutic molecules. For example, kinases play critical roles in regulating cell growth and survival pathways implicated in cancer. Thus, predicting kinase–ligand binding affinity facilitates the effective screening of kinase inhibitors, which can selectively block these pathways. The screened kinase inhibitors are promising candidates for targeted cancer therapy. Predicting protein–ligand binding affinity can be formulated as a regression task: $f:(\mathcal{P},\mathcal{M})\rightarrow\mathbb{R}$, where $\mathcal{P}$ denotes the protein and $\mathcal{M}$ denotes the ligand molecule.
PLI Domain Models. Some domain-specific methods, such as MONN (Li et al., 2020b), do not rely on structural information. Instead, they predict atom–residue noncovalent interaction matrices as intermediate mechanistic signals to assist binding affinity prediction. To mitigate the false positive issue, KDBNet (Luo et al., 2023b) applied an uncertainty recalibration technique to refine the uncertainty estimates.
@@
Table 2. Comparisons across model architectures under different training strategies. The first line for each model (e.g., EGNN) denotes a randomly initialized protein encoder trained purely with downstream task supervision. The subsequent lines (e.g., w/MLM) represent pretrained encoders with frozen weights, where only the task-specific head is fine-tuned. Results of encoder finetuning are in Appendix D.3. Mutation effect prediction (MTP) is only applicable to masked language modeling.
Model Specialized Tasks General Tasks
PCS PROTACs PLI PFA MTP
@@
D-Trans 97.60 $\pm$ 0.08 61.42 $\pm$ 1.30 88.28 $\pm$ 0.79 22.82 $\pm$ 2.95 80.08 $\pm$ 1.08 86.58 $\pm$ 0.28 0.416 $\pm$ 1.028 60.92 $\pm$ 0.22 19.57 $\pm$ 0.80 11.16 $\pm$ 0.67
w/MLM 96.53 $\pm$ 0.52 20.12 $\pm$ 2.86 84.10 $\pm$ 0.66 11.73 $\pm$ 1.41 77.78 $\pm$ 1.93 85.16 $\pm$ 0.89 0.494 $\pm$ 1.116 55.79 $\pm$ 0.50 17.92 $\pm$ 0.14 9.85 $\pm$ 0.09 -0.09 56.84
w/MVCL 95.23 $\pm$ 0.46 41.26 $\pm$ 10.3 83.50 $\pm$ 0.79 15.85 $\pm$ 1.30 74.16 $\pm$ 1.44 82.32 $\pm$ 0.81 0.550 $\pm$ 2.620 54.21 $\pm$ 0.08 15.87 $\pm$ 0.26 8.00 $\pm$ 0.20
w/PFP 96.40 $\pm$ 0.59 39.56 $\pm$ 5.55 85.51 $\pm$ 0.42 12.29 $\pm$ 1.26 74.76 $\pm$ 1.08 82.34 $\pm$ 0.09 0.445 $\pm$ 1.064 56.46 $\pm$ 0.01 17.46 $\pm$ 0.18 9.46 $\pm$ 0.12
ESM-2 97.23 $\pm$ 0.06 41.22 $\pm$ 0.25 86.34 $\pm$ 0.12 6.16 $\pm$ 0.07 78.46 $\pm$ 3.03 85.74 $\pm$ 3.47 0.491 $\pm$ 4.823 53.25 $\pm$ 1.26 49.79 $\pm$ 0.03 43.44 $\pm$ 0.17 43.05 73.48
ESM-C 95.76 $\pm$ 0.01 38.28 $\pm$ 0.08 88.08 $\pm$ 0.00 8.06 $\pm$ 0.08 80.47 $\pm$ 0.90 86.74 $\pm$ 0.37 0.461 $\pm$ 0.037 53.41 $\pm$ 0.87 38.27 $\pm$ 0.16 30.31 $\pm$ 0.12 42.51 73.24
SaProt 96.16 $\pm$ 0.16 42.91 $\pm$ 0.27 89.18 $\pm$ 0.06 13.48 $\pm$ 0.03 73.44 $\pm$ 1.32 80.29 $\pm$ 0.70 0.473 $\pm$ 0.004 51.66 $\pm$ 0.67 33.06 $\pm$ 0.09 25.21 $\pm$ 0.11 49.22 76.89
General Application 4: Protein Function Annotation Prediction (PFA).
Application Description of PFA. To systematically represent protein functions, the Gene Ontology (GO) annotation framework was developed (Ashburner et al., 2000), providing structured annotations across Molecular Function (MF), Cellular Component (CC), and Biological Process (BP). E.g., in the Molecular Function (MF) aspect, a protein may be annotated with specific biochemical activities, such as immune Response or signal transduction. Among the known protein sequences, fewer than 1% have experimentally validated functional annotations. Predicting the GO terms of proteins helps researchers better understand protein functions and potentially guide the discovery and design of proteins with desired functional properties.
PFA Domain Models. Protein function prediction is inherently a multi-label classification task. Early methods relied on sequence alignment or direct modeling of the sequence. More recent approaches have progressively integrated additional modalities. E.g., DeepFRI (Gligorijević et al., 2021) combines protein structure and pre-trained sequence embeddings using a GCN, DPFunc (Wang et al., 2025) further incorporates domain-level information, and the PLM-based DeepGO-SE (Kulmanov et al., 2024).
General Application 5: Mutation Effect Prediction for Protein Optimization (MTP).
Application Description of MTP. Protein mutations refer to the substitution, insertion, or deletion of amino acids within protein sequences. As illustrated in Fig. 3, mutations are frequently associated with changes in functional properties such as stability, binding affinity, and pathogenicity. Computational models can explore the protein fitness landscape to map sequences or structures to functional properties, which can enable protein optimization with mutations. For example, mutations in antibodies or protein complexes can be systematically explored to discover variants with enhanced binding affinity (Cai et al., 2024). MTP can be formulated as estimating changes in target properties (e.g., stability changes measured by $\Delta\Delta G$) resulting from protein mutations. Due to limited experimental annotations, mutation effect prediction typically relies on a zero-shot learning setting.
MTP Domain Models. Recently, protein language models (PLMs) have shown promise in mutation effect prediction due to their ability to capture patterns learned from evolutionary data. PLMs tend to assign higher probabilities to mutations that are consistent with evolutionary patterns, and these mutations are more likely to yield beneficial functional effects. Details of the zero-shot mutation effective prediction with PLMs can be found in Appendix F.5.
5. Experiments
In this section, we conduct empirical studies to compare domain models and various pretraining frameworks and strategies on the downstream applications described in Sec. 4.
5.1. Experimental Setup
5.1.1. Pretraining Setup
We describe the pretraining setting here.
Pretraining Dataset. For pre-training, we collect 542,378 protein structures from AlphaFold Protein Structure Database (AFDB) and obtain the corresponding amino acid sequences and protein family labels from UniProt (Varadi et al., 2022; Consortium, 2019). This curated dataset spans diverse and non-redundant protein structures, enabling a comprehensive evaluation of different training strategies. ESM-2 and SaProt are trained on protein sequences from UniRef (Consortium, 2019), while ESM-C further incorporates sequences from MGnify (Richardson et al., 2023) and the JGI (Nordberg et al., 2014). Additional dataset details are provided in Appendix G.2.
Model Setups. We use fixed random seeds 42,128,256,512,1024{42,128,256,512,1024} for reproducibility, and report mean $\pm$ standard deviation over five runs. All models are trained with the Adam optimizer. The learning rate follows a linear warm-up over 10% of training steps, then cosine annealing to 1e-6, with a peak of 1e-4. More training details are provided in Appendix F.
5.1.2. Downstream Task Training Setup
In this subsection, we introduce the setup of training pretrained models and domain models for downstream application tasks.
Downstream Application Dataset. Each downstream task uses the same dataset for both training-from-scratch and fine-tuning to ensure fair comparison. All datasets are collected from peer-reviewed sources and standardized through quality control and formatting, as detailed in Appendix F. Specifically, PROTACs data are from PROTAC-DB (Ge et al., 2025), cleavage data from MEROPS (Rawlings et al., 2014), protein–ligand interactions from KDBNet, function prediction from DeepFRI (Gligorijević et al., 2021) (with structural data from ProteinWorkshop (Jamasb et al., 2024)), and mutation prediction from ProteinGym (Notin et al., 2023).
Molecular Encoders in Downstream Applications. The applications of PLI and PROTACs involve various chemical molecules in the interactions with proteins, requiring a molecular encoder. To maintain consistency and comparability across interaction tasks, molecular components, drugs in PLI, warheads, linkers, and E3-ligands in PROTACs are encoded using randomly initialized GVP encoders and updated under all training settings. For domain-specific models, molecular encoders follow their original implementations.
Downstream Task Training Strategies. All domain-specific models are trained from scratch. Pretrained models serve as protein encoders under the following training settings:
  • Training from scratch: All parameters are randomly initialized and trained end-to-end.
  • Freeze-encoder training: Encoders are initialized with pretrained weights and kept frozen, while only task-specific heads and molecular encoders are updated.
  • Multi-stage fine-tuning: The protein encoder initialized with pretrained weights is first frozen to train the task-specific heads and molecular encoders, followed by fine-tuning the entire model using a smaller learning rate.
Evaluation. We evaluate all methods on the five downstream applications in Protap using task-standard metrics which are shown in Table. 1(a). Detailed criteria for selecting these metrics are provided in Appendix F.
(a) Protein Cleavage Site
(b) Proteolysis-Targeting Chimeras
(c) Protein–Ligand Interactions
(d) Protein Function Annotation
A short description of the image A long description of the image content, detailing all important visual information for a reader using screen-reading software.
Figure 4. Performance comparison between general architecture and domain-specific models across downstream applications.
Table 3. Summary of best results in different pretrain objectives under each fine-tuning strategy.
Model Fine-tuning strategy PCS PROTACs PLI
C14.005 PROTACDB Davis